p=0.64

a=0.05

n=180


what is the point estimate of p?


what is the interval estimate of p?


1
Expert's answer
2022-05-29T17:51:48-0400

a)

The point estimate of "p" is "p=0.64"


b)

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}}, \hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(0.64-1.96\sqrt{\dfrac{0.64(1-0.64)}{180}},"

"0.64+1.96\sqrt{\dfrac{0.64(1-0.64)}{180}})"

"=(0.57, 0.71)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.57 < p < 0.71," which indicates that we are 

95% confident that the true population proportion "p"  is contained by the interval "(0.57, 0.71)."